I'm stuck on some questions could i give u 2 of them and it might show me how to do the rest ...thank you
8(w-1)-7w
3(2m+4)+2(6-3m)

Answer by stanbon(48568)

8(w-1)-7w = 8w-8-7w = w-8
3(2m+4)+2(6-3m) = 6m+12+12-6m = 24
Cheers,
Stan H.